Given the following data and a bigram language model that uses Laplace smoothing, what is the estimate that the word shark
Answer #1

Bigram LM with Laplace smoothing:

P(wn | wn-1) = C( wn-1wn)+1/C(wn) +V

V is vocabulary size for a bigram model

P(shark | fish) = C(fish,shark)+1/C(fish)+V

C(fish,shark)= 0

C(fish) =4

V= 7 (unique words are: a,shark,is, fish,sharky,are,too)

P(shark | fish) = 0+1/4+7 = 1/11=0.091

Option E is correct.